# EndoGeDE: Generalizable Monocular Depth Estimation with Mixture of Low-Rank Experts for Diverse Endoscopic Scenes

### v2

@@ 3.4.3 Third Stage: Multiple Transformation Alignments
4 Experiments and Results
4.1 Implementation Details and Metrics
@@
4.2 Datasets
SCARED: SCARED(Allan et al., 2021) dataset consists of 35 videos collected from porcine cadavers using a da Vinci Xi surgical robot. Following (Shao et al., 2022), this dataset is divided into 15351, 1705, and 551 frames for training, validation, and test, respectively.
Hamlyn: Hamlyn(Recasens et al., 2021) dataset contains numerous endoscopic videos of diverse robotic surgeries. In the experiments, all 92,672 endoscopic frames from the 21 rectified videos are used for zero-shot depth estimation.
SERV-CT: SERV-CT(Edwards et al., 2022) dataset consists of 8 keyframes collected from two ex-vivo porcine cadavers. In the experiments, all 32 endoscopic frames are used for zero-shot depth estimation.
SimCol: SimCol(Rau et al., 2024) dataset contains over 36,000 colonoscopic images based on the simulation in Unity3D. Following official website, the dataset is split into 28,776 and 9,009 frames for training and test, respectively.
C3VD: C3VD(Bobrow et al., 2023) dataset consists of 22 short video sequences collected from high-fidelity colon models by a clinical colonoscope. Following previous works (Paruchuri et al., 2024), 2889 frames from 8 videos are selected for zero-shot depth estimation.
EndoMapper: EndoMapper(Azagra et al., 2023) dataset consists of realistic colonoscopic sequences collected from regular medical practice. For sim-to-real evaluation, 2 frames are selected for the display of zero-shot depth estimation.
Table 1: Comparison Results of Self-supervised Depth Estimation on SCARED. The BEST results are in BOLD. $\dagger$ denotes that the foundation model is finetuned based on (Cui et al., 2024b).
Methods Backbones Year $Rel_{Abs}\downarrow$ $Rel_{Sq}\downarrow$ $RMSE\downarrow$ $RMSE_{Log}\downarrow$ $\delta\uparrow$ TP/M
AF-SfMLearner (Shao et al., 2022) CNN 2022 0.059 0.435 4.925 0.082 0.974 14.8
MonoViT (Zhao et al., 2022) ViT 2022 0.062 0.470 5.042 0.082 0.976 27.9
LiteMono (Zhang et al., 2023a) CA 2023 0.057 0.453 4.967 0.079 0.975 3.1
Surgical-DINO (Cui et al., 2024a) DN 2024 0.059 0.427 4.904 0.081 0.974 0.1
Yang et al.(Yang et al., 2024c) CA 2024 0.062 0.558 5.585 0.090 0.962 2.0
@@ 4.2 Datasets
IID-SfMLearner(Li et al., 2024) CNN 2024 0.058 0.435 4.820 0.080 0.969 14.8
DVSMono (Zhou et al., 2024) MV 2024 0.055 0.410 4.797 0.078 0.975 27.0
DARES (Zeinoddin et al., 2024) DA 2024 0.052 0.356 4.483 0.073 0.980 2.9
EndoDAC(Cui et al., 2024b) DA 2024 0.052 0.362 4.464 0.072 0.979 1.6
SfM-Diffusion (Li et al., 2025) SD 2025 0.049 0.366 4.305 0.078 0.975 12.3
MonoPCC(Wang et al., 2025) MV 2025 0.051 0.349 4.488 0.072 0.983 27.0
[…]
CA: CNN+Attention, DN: DINO, MV: MonoViT, DA: Depth Anything, SD: Stable Diffusion.
4.3 Depth Estimation Results
4.3.1 Results on Realistic Data
The proposed method is evaluated on three different realistic datasets, SCARED, Hamlyn and SERV-CT. Meanwhile, our method is compared with state-of-the-art works including methods for endoscopic scenes (MonoPCC(Wang et al., 2025), EndoDAC(Cui et al., 2024b), DVSMono(Zhou et al., 2024), DARES(Zeinoddin et al., 2024), SfM-Diffusion(Li et al., 2025), IID-SfM(Li et al., 2024), Surgical-DINO(Cui et al., 2024a), AF-SfM(Shao et al., 2022) and method from (Yang et al., 2024c)) and natural scenes (MonoViT(Zhao et al., 2022), LiteMono(Zhang et al., 2023a) and Depth Anything(Yang et al., 2024a, b)). All frameworks are trained and tested on SCARED dataset at first, results of which are shown in Table. 4.2 and Fig. 6. To evaluate the generalization of the model, all models trained on SCARED dataset are utilized to perform zero-shot depth estimation on Hamlyn and SERV-CT dataset, and the experimental results are shown in Table. 2 and Fig. 7.
Table 2: Comparison Results of Zero-shot Depth Estimation on Hamlyn and SERV-CT Dataset. The BEST results are in BOLD.
@@ 4.3.1 Results on Realistic Data
AF-SfMLearner(Shao et al., 2022) 0.168 4.440 13.870 0.204 0.770 0.105 1.772 11.590 0.134 0.889
MonoViT (Zhao et al., 2022) 0.193 10.512 18.028 0.220 0.769 0.103 1.566 11.482 0.136 0.895
LiteMono (Zhang et al., 2023a) 0.179 6.366 15.196 0.216 0.754 0.124 2.314 13.156 0.175 0.820
Depth Anything(Yang et al., 2024a) 0.154 3.616 12.733 0.189 0.784 0.082 1.122 9.409 0.110 0.929
Depth Anything v2(Yang et al., 2024b) 0.182 4.994 15.067 0.219 0.740 0.092 1.328 10.407 0.120 0.918
IID-SfMLearner(Li et al., 2024) 0.171 4.526 14.066 0.206 0.767 0.112 1.956 12.193 0.138 0.878
DVSMono(Zhou et al., 2024) 0.143 2.956 11.905 0.181 0.796 0.095 1.412 10.811 0.127 0.907
DARES(Zeinoddin et al., 2024) 0.164 4.456 13.625 0.201 0.783 0.084 1.179 9.932 0.113 0.930
EndoDAC(Shao et al., 2022) 0.156 3.854 12.936 0.193 0.791 0.084 1.223 9.683 0.113 0.922
MonoPCC(Wang et al., 2025) 0.158 3.889 13.205 0.194 0.782 0.091 1.265 10.123 0.117 0.915
[…]
Figure 6: Qualitative results of depth estimation on SCARED dataset. The error maps denotes the relative error distribution in the area marked by the green box.
4.3.2 Results on Simulated Data and Sim-to-Real
To further demonstrate the performance of our method, the proposed framework is trained and evaluated on simulated datasets. Firstly, our method is compared with existing methods, which are with public training codes, on SimCol dataset. Meanwhile, the models trained on SimCol dataset are compared with previous works for zero-shot depth estimation on C3VD dataset. Furthermore, the proposed model with the pretrained weights from SimCol is also utilized to perform depth estimation on samples from EndoMapper, which can display the sim-to-real performance of our work. The quantitative results are shown in Table 3, while qualitative results including sim-to-real depth estimation are shown in Fig. 8. According to the experimental results, the outstanding performance and generalization of the proposed method are further proved.
Table 3: Comparison Results of Depth Estimation on SimCol and C3VD Dataset. The BEST results are in BOLD.
@@ 4.3.2 Results on Simulated Data and Sim-to-Real
AF-SfMLearner(Shao et al., 2022) 0.086 0.080 0.426 0.109 0.950 0.142 1.495 7.897 0.188 0.804
IID-SfMLearner(Li et al., 2024) 0.080 0.069 0.429 0.108 0.959 0.140 1.484 7.733 0.186 0.810
EndoDAC(Cui et al., 2024b) 0.107 0.110 0.530 0.133 0.924 0.103 0.652 5.306 0.127 0.918
[…]
@@
Table 4: Ablations on Parameter-efficient Finetuning Strategy. The best results are in red.
@@
HydraLoRA(Tian et al., 2024) 0.052 / 0.098 / 0.155 0.361 / 1.675 / 4.066 4.525 / 10.931 / 12.986 0.073 / 0.125 / 0.191 0.982 / 0.904 / 0.796
AdaLoRA(Zhang et al., 2023b) 0.054 / 0.076 / 0.144 0.438 / 0.985 / 3.057 4.949 / 8.992 / 12.075 0.078 / 0.103 / 0.179 0.978 / 0.933 / 0.801
DV-LoRA(Cui et al., 2024b) 0.050 / 0.086 / 0.151 0.334 / 1.172 / 3.639 4.342 / 9.403 / 12.469 0.071 / 0.108 / 0.186 0.983 / 0.929 / 0.798
LoRA(Hu et al., 2022) 0.049 / 0.081 / 0.150 0.336 / 1.065 / 3.681 4.381 / 9.035 / 12.508 0.070 / 0.103 / 0.185 0.983 / 0.945 / 0.800
4.4 Ablation Studies
4.4.1 Ablations on Parameter-efficient Finetuning
Ablation studies are performed to demonstrate the effect of the proposed parameter-efficient finetuning based on block-wise mixture of low-rank experts (BW-MoLE), as Table 4.3.2 shows. Firstly, the allocation of the experts are replaced with group-wise allocation (GW-MoLE(Gao et al., 2025)) and average allocation. Note that the deeper group is also allocated with more experts in GW-MoLE according to Fig. 7. The numbers of experts are set to 2, 4, 6 and 8 for each group consisting of four Transformer blocks for group-wise allocation, while the top-ks in all blocks are equal to 2. Moreover, the number of experts is set to x with the top-k setting, represented as MoLE(k/x), for average allocation. Furthermore, the finetuning of each linear layer is demonstrated based on ablation studies. At last, the finetuning based on the proposed method is also compared with the PEFT with multiple kinds of LoRA.
4.4.2 Ablations on Self-supervised Learning
Ablation studies are performed to demonstrate the effect of the proposed self-supervised training strategy, as Table 5 shows. For our training strategy, the experiments are implemented to prove the effect of the loss from augmented image, Retinex algorithm and illumination-free alignment. Moreover, the proposed model is also trained based on the existing training pipeline AF-SfMLearner (Shao et al., 2022) and IID-SfMLearner (Liu et al., 2020) for comparison.
Table 5: Ablation studies on Self-supervised Learning (SSL) pipeline. The best results are in red.
@@ 4.4.2 Ablations on Self-supervised Learning
AF-SfMLearner(Shao et al., 2022) 0.049 / 0.083 / 0.142 0.339 / 1.071 / 3.138 4.390 / 9.086 / 11.842 0.070 / 0.107 / 0.180 0.983 / 0.932 / 0.805
IID-SfMLearner(Li et al., 2024) 0.058 / 0.092 / 0.148 0.448 / 1.342 / 3.302 4.997 / 9.924 / 12.372 0.081 / 0.118 / 0.184 0.976 / 0.913 / 0.794
4.5 3D Reconstruction Results
3D reconstruction is a significant application of depth estimation for endoscopy. Therefore, we also perform 3D scene reconstruction based on the estimated depth maps using Open3D  [^3] https://www.open3d.org/. As Fig. 9 shows, the reconstructed scenes based on the depth maps provided by our method is with more clear and accuracy texture and appearance, compared with results from (Cui et al., 2024b) and (Wang et al., 2025).
@@
4.6 Ego-motion Estimation Results
Ego-motion estimation, as one of important part in the self-supervised framework, is also evaluated with experiments. Four sequences from the test split of SCARED dataset are selected to perform ego-motion estimation. The accuracy of ego-motion estimation is evaluated and compared based on Absolute Trajectory Error (ATE), following previous works. Based on experimental results in Table 6 and Fig. 10, the proposed method can perform the most accurate ego-motion estimation, compared with multiple existing methods including MonoPCC(Wang et al., 2025), EndoDAC(Cui et al., 2024b), IID-SfM(Li et al., 2024) and AF-SfM(Shao et al., 2022).
Table 6: Ego-motion estimation results on SCARED dataset
@@ 4.6 Ego-motion Estimation Results
AF-SfMLearner(Shao et al., 2022) 0.0941 0.0841 0.0742 0.0682
Depth Anything(Yang et al., 2024a) 0.0932 0.0815 0.0794 0.0696
IID-SfMLearner(Li et al., 2024) 0.0951 0.0851 0.0764 0.0682
EndoDAC(Cui et al., 2024b) 0.0936 0.0832 0.0776 0.0687
MonoPCC(Wang et al., 2025) 0.1040 0.0842 0.0781 0.0939
[…]
5 Discussion
5.1 Outstanding Accuracy and Generalization
@@
5.2 Generalization on Diverse Scenes
@@
5.3 Generalization on Diverse Illumination
@@
5.4 Limitations & Future Works
Although the proposed method achieves state-of-the-art performance compared with existing methods on monocular depth estimation in diverse endoscopic scenes, The inference speed of our method is lower than previous works while the speed still achieves about 30fps which could meet the real-time requirement, shown in Table.7. Although the number of training parameters remains low in the proposed framework, the step-by-step training strategy and the usage of foundation model still need lots of computational resources. In the future, more efficient training pipeline and foundation model could be developed for endoscopic depth estimation. Meanwhile, multimodal visual features such as semantic feature could boost the performance of depth estimation in endoscopy.
Table 7: Comparison of Inference Speed
@@ 5.4 Limitations & Future Works
Depth Anything(Yang et al., 2024a, b) 5.0 3.0
IID-SfMLearner(Li et al., 2024) 2.0 1.0
DVSMono(Zhou et al., 2024) 12.7 10.5
EndoDAC(Cui et al., 2024b) 5.7 4.0
MonoPCC(Wang et al., 2025) 12.6 10.5
[…]
6 Conclusion

### v3

@@ III-E Stage III: Multiple Transformation Alignments
IV Experiments and Results
IV-A Implementation Details and Metrics
@@
TABLE I: Comparison Results of Self-supervised Depth Estimation on SCARED. The BEST results are in BOLD. $\dagger$ denotes that the foundation model is finetuned based on [15]. TP denotes the number of trainable parameters.
Methods Year $Rel_{Abs}\downarrow$ $Rel_{Sq}\downarrow$ $RMSE\downarrow$ $RMSE_{Log}\downarrow$ $\delta\uparrow$ TP/M
AF-SfMLearner [12] 2022 0.059 0.435 4.925 0.082 0.974 14.8
MonoViT [10] 2022 0.062 0.470 5.042 0.082 0.976 27.9
LiteMono [11] 2023 0.057 0.453 4.967 0.079 0.975 3.1
Surgical-DINO [14] 2024 0.059 0.427 4.904 0.081 0.974 0.1
Yang et al.[8] 2024 0.062 0.558 5.585 0.090 0.962 2.0
@@
IID-SfMLearner[13] 2024 0.058 0.435 4.820 0.080 0.969 14.8
DVSMono [31] 2024 0.055 0.410 4.797 0.078 0.975 27.0
DARES [16] 2024 0.052 0.356 4.483 0.073 0.980 2.9
EndoDAC[15] 2024 0.052 0.362 4.464 0.072 0.979 1.6
SfM-Diffusion [40] 2025 0.049 0.366 4.305 0.078 0.975 12.3
MonoPCC[32] 2025 0.051 0.349 4.488 0.072 0.983 27.0
EndoGMDE Ours
0.047
↓4.1%
0.307
↓12%
4.206
↓2.3%
0.067
↓6.9%
0.985
↑0.2%
3.7
$\Delta$
TABLE II: Comparison Results of Zero-shot Depth Estimation on Hamlyn and SERV-CT Dataset. The BEST results are in BOLD.
@@
AF-SfMLearner[12] 0.168 4.440 13.870 0.204 0.770 0.105 1.772 11.590 0.134 0.889
MonoViT [10] 0.193 10.512 18.028 0.220 0.769 0.103 1.566 11.482 0.136 0.895
LiteMono [11] 0.179 6.366 15.196 0.216 0.754 0.124 2.314 13.156 0.175 0.820
Depth Anything[34] 0.154 3.616 12.733 0.189 0.784 0.082 1.122 9.409 0.110 0.929
Depth Anything v2[35] 0.182 4.994 15.067 0.219 0.740 0.092 1.328 10.407 0.120 0.918
IID-SfMLearner[13] 0.171 4.526 14.066 0.206 0.767 0.112 1.956 12.193 0.138 0.878
DVSMono[31] 0.143 3.056 11.905 0.181 0.796 0.095 1.412 10.811 0.127 0.907
DARES[16] 0.164 4.456 13.625 0.201 0.783 0.084 1.179 9.932 0.113 0.930
EndoDAC[12] 0.156 3.854 12.936 0.193 0.791 0.084 1.223 9.683 0.113 0.922
MonoPCC[32] 0.158 3.889 13.205 0.194 0.782 0.091 1.265 10.123 0.117 0.915
EndoGMDE
0.140
↓2.1%
3.092
↑1.2%
11.749
↓1.3%
0.175
↓3.3%
0.808
↑1.5%
0.069
↓16%
0.307
↓35%
7.597
↓19%
0.088
↓20%
0.963
↑3.5%
$\Delta$
IV-B Datasets
SCARED: SCARED[19] dataset consists of 35 videos collected from porcine cadavers using a da Vinci Xi surgical robot. Following [12], this dataset is divided into 15351, 1705, and 551 frames for training, validation, and test, respectively.
Hamlyn: Hamlyn[22] dataset contains numerous endoscopic videos of diverse robotic surgeries. In the experiments, all 92,672 endoscopic frames from the 21 rectified videos are used for zero-shot depth estimation.
SERV-CT: SERV-CT[23] dataset consists of 8 keyframes collected from two ex-vivo porcine cadavers. In the experiments, all 32 endoscopic frames are used for zero-shot depth estimation.
SimCol: SimCol[20] dataset contains over 36,000 colonoscopic images based on the simulation in Unity3D. Following official website, the dataset is split into 28,776 and 9,009 frames for training and test, respectively.
C3VD: C3VD[21] dataset consists of 22 short video sequences collected from high-fidelity colon models by a clinical colonoscope. Following previous works [41], 2889 frames from 8 videos are selected for zero-shot depth estimation.
EndoMapper: EndoMapper[24] dataset consists of realistic colonoscopic sequences collected from regular medical practice. For sim-to-real evaluation, 2 frames are selected for the display of zero-shot depth estimation.
IV-C Depth Estimation Results
Figure 6: Qualitative results of depth estimation on SCARED dataset. The error maps denotes the relative error distribution in the area marked by the green box. The first image at the row of error maps is the ground truth of depth map in corresponding area.
TABLE III: Comparison Results of Depth Estimation on SimCol and C3VD Dataset. The BEST results are in BOLD.
@@ IV-C Depth Estimation Results
AF-SfMLearner[12] 0.086 0.080 0.426 0.109 0.950 0.142 1.495 7.897 0.188 0.804
IID-SfMLearner[13] 0.080 0.069 0.429 0.108 0.959 0.140 1.484 7.733 0.186 0.810
EndoDAC[15] 0.107 0.110 0.530 0.133 0.924 0.103 0.652 5.306 0.127 0.918
EndoGMDE
0.078
↓2.5%
0.056
↓19%
0.421
↓1.2%
0.109
↑0.9%
0.959
0%
0.096
↓6.8%
0.635
↓2.6%
5.241
↓1.2%
0.120
↓5.5%
0.923
↑0.5%
$\Delta$
IV-C1 Results on Realistic Data
The proposed method is evaluated on three different realistic datasets, SCARED, Hamlyn and SERV-CT. Meanwhile, our method is compared with state-of-the-art works including methods for endoscopic scenes (MonoPCC[32], EndoDAC[15], DVSMono[31], DARES[16], SfM-Diffusion[40], IID-SfM[13], Surgical-DINO[14], AF-SfM[12] and method from [8]) and natural scenes (MonoViT[10], LiteMono[11] and Depth Anything[34, 35]). All frameworks are trained and tested on SCARED dataset at first, results of which are shown in Table. I and Fig. 6. To evaluate the generalizability of the model, all frameworks trained on SCARED dataset are utilized to perform zero-shot depth estimation on Hamlyn and SERV-CT dataset, and the experimental results are shown in Table. II and Fig. 7. Note that some works without published codes and weights, including Surgical-DINO[14], Yang’s method[8] and SfM-Diffusion[40], are not compared for zero-shot estimation.
@@ IV-C1 Results on Realistic Data
IV-C2 Results on Simulated Data and Sim-to-Real
To further demonstrate the performance of our method, the proposed framework is trained and evaluated on simulated datasets. Firstly, our method is compared with existing methods, which are with public training codes, on SimCol dataset. Meanwhile, the models trained on SimCol dataset are compared with previous works for zero-shot depth estimation on C3VD dataset. Furthermore, the proposed model with the pretrained weights from SimCol is also utilized to perform depth estimation on samples from EndoMapper, which can display the sim-to-real performance of our work. The quantitative results are shown in Table III, while qualitative results including sim-to-real depth estimation are shown in Fig. 8. According to the experimental results, the outstanding performance and generalization of the proposed method are further proved.
TABLE IV: Ablations on Parameter-efficient Finetuning Strategy. The best results are in red.
@@ IV-C2 Results on Simulated Data and Sim-to-Real
HydraLoRA[43] 0.052 / 0.098 / 0.155 0.361 / 1.675 / 4.066 4.525 / 10.931 / 12.986 0.073 / 0.125 / 0.191 0.982 / 0.904 / 0.796
AdaLoRA[44] 0.054 / 0.076 / 0.144 0.438 / 0.985 / 3.057 4.949 / 8.992 / 12.075 0.078 / 0.103 / 0.179 0.978 / 0.933 / 0.801
DV-LoRA[15] 0.050 / 0.086 / 0.151 0.334 / 1.172 / 3.639 4.342 / 9.403 / 12.469 0.071 / 0.108 / 0.186 0.983 / 0.929 / 0.798
LoRA[17] 0.049 / 0.081 / 0.150 0.336 / 1.065 / 3.681 4.381 / 9.035 / 12.508 0.070 / 0.103 / 0.185 0.983 / 0.945 / 0.800
TABLE V: Ablation studies on Self-supervised Learning (SSL) pipeline. The best results are in red.
@@
AF-SfMLearner[12] 0.049 / 0.083 / 0.142 0.339 / 1.071 / 3.138 4.390 / 9.086 / 11.842 0.070 / 0.107 / 0.180 0.983 / 0.932 / 0.805
IID-SfMLearner[13] 0.058 / 0.092 / 0.148 0.448 / 1.342 / 3.302 4.997 / 9.924 / 12.372 0.081 / 0.118 / 0.184 0.976 / 0.913 / 0.794
IV-D Ablation Studies
IV-D1 Ablations on Parameter-efficient Finetuning
Ablation studies are performed to demonstrate the effect of the proposed parameter-efficient finetuning based on block-wise mixture of low-rank experts (BW-MoLE), as Table IV shows. Firstly, the allocation of the experts are replaced with group-wise allocation (GW-MoLE[42]) and average allocation. Note that the deeper group is also allocated with more experts in GW-MoLE according to Fig. 11. The numbers of experts are set to 2, 4, 6 and 8 for each group consisting of four Transformer blocks for group-wise allocation, while the top-ks in all blocks are equal to 2. Moreover, the number of experts is set to x with the top-k setting, represented as MoLE(k/x), for average allocation. Furthermore, the finetuning of each linear layer is demonstrated based on ablation studies. At last, the finetuning based on the proposed method is also compared with the PEFT with multiple kinds of LoRA.
IV-D2 Ablations on Self-supervised Learning
Ablation studies are performed to demonstrate the effect of the proposed self-supervised training strategy, as Table V shows. For our training strategy, the experiments are implemented to prove the effect of the loss from augmented image, Retinex algorithm and illumination-free alignment. Moreover, the proposed model is also trained based on the existing training pipeline AF-SfMLearner [12] and IID-SfMLearner [45] for comparison.
@@ IV-D2 Ablations on Self-supervised Learning
IV-E 3D Reconstruction Results
3D reconstruction is a significant application of depth estimation for endoscopy. Therefore, we also perform 3D scene reconstruction based on the estimated depth maps using Open3D (https://www.open3d.org/). As Fig. 9 shows, the reconstructed scenes based on the depth maps provided by our method is with more clear and accuracy texture and appearance, compared with results from [15] and [32].
@@
IV-F Ego-motion Estimation Results
Ego-motion estimation, as one of important part in the self-supervised framework, is also evaluated with experiments. Four sequences from the test split of SCARED dataset are selected to perform ego-motion estimation. The accuracy of ego-motion estimation is evaluated and compared based on Absolute Trajectory Error (ATE), following previous works. Based on experimental results in Table VI and Fig. 10, the proposed method can perform the most accurate ego-motion estimation, compared with multiple existing methods including MonoPCC[32], EndoDAC[15], IID-SfM[13] and AF-SfM[12].
TABLE VI: Ego-motion estimation results on SCARED dataset
@@ IV-F Ego-motion Estimation Results
AF-SfMLearner[12] 0.0941 0.0841 0.0742 0.0682
Depth Anything[34] 0.0932 0.0815 0.0794 0.0696
IID-SfMLearner[13] 0.0951 0.0851 0.0764 0.0682
EndoDAC[15] 0.0936 0.0832 0.0776 0.0687
MonoPCC[32] 0.1040 0.0842 0.0781 0.0939
Ours 0.0927 0.0814 0.0754 0.0678
V Discussion
V-A Outstanding Accuracy and Generalization
@@
V-B Generalization on Diverse Scenes
@@
V-C Generalization on Diverse Illumination
@@
TABLE VII: Comparison of Inference Speed
@@
Depth Anything[34, 35] 5.0 3.0
IID-SfMLearner[13] 2.0 1.0
DVSMono[31] 12.7 10.5
EndoDAC[15] 5.7 4.0
MonoPCC[32] 12.6 10.5
Ours 34.9($\sim$30fps) 32.5(>>30fps)
V-D Limitations & Future Works
Although the proposed method achieves state-of-the-art performance compared with existing methods on monocular depth estimation in diverse endoscopic scenes, The inference speed of our method is lower than previous works while the speed still achieves about 30fps which could meet the real-time requirement, shown in Table.VII. Although the number of training parameters remains low in the proposed framework, the step-by-step training strategy and the usage of foundation model still need lots of computational resources. In the future, more efficient training pipeline and foundation model could be developed for endoscopic depth estimation. Meanwhile, multimodal visual features such as semantic feature could boost the performance of depth estimation in endoscopy.
VI Conclusion